Suppose you have 4 combination locks, each with three spinners with digits 0-9. (so there are 1000 combinations 000-999) what is the probability that at least two of the locks have the same combination?
Answer #1

probability that at least two of the locks have the same combination =1-P(all 4 locks have different

combinations) =1-(1000/1000)*(999/1000)*(998/1000)*(997/1000)=0.005989